A coordinate system (in meters) is constructed on the surface of a pool table, and three objects are placed on the table as follows: a m1 = 1.8-kg object at the origin of the coordinate system, a m2 = 2.9-kg object at

(0, 2.0),

and a m3 = 4.3-kg object at

(4.0, 0).

Find the resultant gravitational force exerted by the other two objects on the object at the origin.

magnitude     N
direction     ° above the +x-axis
Answer #1

y 6 M2=2.8 kg t Fiz My = 4.3kg F12 = GM, M2 72 M=1.8kg (4.00) 6.67x10x 1.882.9 - 8-7410N g F3 = GMIMI 6.67X10 X 1.8x4.3 = 3.F = F1 = 3-22x10h à +8 17x10 9 √(3 +227 ²+ (8.712 x 18 = √86-0584 x 10th .6564 x 10 Fresultant = 9.29 x 10 N